# Quantum droplets in two-dimensional Bose mixtures at finite temperature

### v1

@@ Acknowledgements.
S.P. and G.S. acknowledge support from the Italian MUR under the PRIN2022 project “Hybrid algorithms for quantum simulators” No. 2022H77XB7, and from CINECA for the availability of high performance computing resources and support under the ISCRA initiative. S.P. also acknowledges support from the PNRR MUR Project No. PE0000023-NQSTI and from the EuroHPC Joint Undertaking, for awarding access to the EuroHPC supercomputer LUMI, hosted by CSC (Finland) and the LUMI consortium through a EuroHPC Benchmark Access call. S.G. acknowledges support from ICSC – Centro Nazionale di Ricerca in HPC, Big Data and Quantum Computing, funded by the European Union under NextGenerationEU. Views and opinions expressed are, however, those of the author(s) only and do not necessarily reflect those of the European Union or The European Research Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. G.S. and S.G. also acknowledge funding from the Provincia Autonoma di Trento.

### v2

@@ Acknowledgements.
Acknowledgments—S.P. and G.S. acknowledge support from the Italian Ministry of University and Research under the PRIN2022 project “Hybrid algorithms for quantum simulators” No. 2022H77XB7, and from CINECA for the availability of high performance computing resources and support under the ISCRA initiative. S.P. also acknowledges support from the PNRR MUR Project No. PE0000023-NQSTI and from the EuroHPC Joint Undertaking, for awarding access to the EuroHPC supercomputer LUMI, hosted by CSC (Finland) and the LUMI consortium through a EuroHPC Benchmark Access call. G.S. and S.G. acknowledge funding from the Provincia Autonoma di Trento. S.G. also acknowledges support from ICSC – Centro Nazionale di Ricerca in HPC, Big Data and Quantum Computing, funded by the European Union under NextGenerationEU. Views and opinions expressed are, however, those of the author(s) only and do not necessarily reflect those of the European Union or The European Research Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.